# The lattice packing problem in dimension 9 by Voronoi’s algorithm

### v1

@@ 1. Introduction
The sphere packing problem is a classical and notoriously hard problem. It asks how to pack equal $n$-dimensional balls in $\mathbb{R}^{n}$ such that their density is maximized. The sphere packing problem has only been solved in dimensions $1,2,3,8$ and $24$, the last two of which only recently [Via17, CKM+17]. Johannes Kepler conjectured already in the 17th century that in dimension $3$ no packing could be denser than the face-centered cubic or hexagonal close packings. A testament of the hardness of this problem is given by the fact that Kepler’s conjecture stayed unproven for almost $400$ years. Until $1998$, when Thomas Hales announced a computational proof to the sphere packing problem in dimension $3$ [Hal05] which was later formalized with proof checking software in [HAB+17].
@@
And indeed, this restriction allows to solve the problem in all dimensions $1$ to $8$ and $24$. Contrary to Kepler’s conjecture the lattice packing problem in dimension $3$ was already solved by Gauss in $1831$ [Gau31]. Dimensions $4$ and $5$ followed in $1877$ by Korkine and Zolotarev’s reduction theory [KZ77], and in $1935$ Blichfeldt used the same theory to determine the optimal lattice packings in dimensions $6,7$ and $8$ [Bli35]. The famous Leech lattice was shown to be optimal in dimension $24$ by Cohn and Kumar [CK09] using linear programming bounds, a technique which also played an important role in solving the general sphere packing problem in dimensions $8$ and $24$. The optimality in all these dimensions was shown by strong theoretical arguments and limited computations. However, the lattice packing problem in dimensions $9$ and beyond (except for $24$) so far seems to evade any of the theoretical approaches that worked for the other dimensions. In particular, all known density upper bounds are relatively far from the best known (lattice) packings in these dimensions.
Almost $100$ years before the computational proof of the Kepler conjecture, Voronoi, ahead of his time, already gave a computational approach to solve the lattice packing problem in any fixed dimension [Vor08]. Voronoi showed that any optimal lattice packing must be a perfect lattice, and that there are only a finite number of non-similar perfect lattices in any given dimension. Furthermore, he presented an algorithm that provably enumerates all perfect lattices, thereby giving a computational way to solve the lattice packing problem.
@@
2 1 Lagrange, 1770 [dL70]
3 1 Gauss, 1831 [Gau31]
4 2 Korkine & Zolotarev, 1877 [KZ77]
5 3 Korkine & Zolotarev, 1877 [KZ77]
6 7 Barnes, 1957 [Bar57]
7 33 Jaquet-Chiffelle, 1993 [JC93]
@@
9 $\geq$ 500 000 Vallentin, 2005 [DSSV07]
$\geq$ 23 000 000 van Woerden, 2018 [vW18]
@@ 9 $\geq$ 500 000 Vallentin, 2005 [DSSV07]
So far, this computational approach to solve the lattice packing problem, has always fallen behind the theoretical results, mainly due to its high complexity. For example, the enumeration of perfect lattices in dimension $6$ was done by Barnes in $1957$ [Bar57], $22$ years after Blichfeldt’s proof. And dimensions $7$ and $8$ followed, only in $1993$ by Jaquet [JC93] and in $2007$ by Dutour Sikirić, Schürmann and Vallentin [DSSV07] respectively. Dimension $9$ was also attempted, and in [DSSV07] and [AA13] about 500 000 perfect lattices were found. In [vW18] some additional algorithmic improvements lead to a total of 23 000 000 perfect lattices to be found. Unfortunately, even with such a large number there was still no end in sight during the enumeration, and provably finding all cases requires a significant amount of extra computation, making it unclear if dimension $9$ would be feasible.
@@
Furthermore, additional data obtained by our computation allows us to prove precisely how many shortest vectors a lattice in dimension $9$ can have.
@@
We aim to make the data and implementations used in this work public. The full classification of perfect forms in dimension $9$ is available split into 64 binary netCDF files at [DSvW25a]. Their total size is roughly 150GB. We wrote a utility to easily access this data once downloaded [^1] Available at https://github.com/MathieuDutSik/AccessPerfectDim9. The dual description under symmetry program has been developed publicly at [DSvW25b]. Our parallel implementation of Voronoi’s algorithm, including the canonical form implementation, will be made public in due course.
@@ Acknowledgements
The authors would like to thank Achill Schürmann and Jacques Martinet for insightful discussions and suggestions that were beneficial to this work.
@@ 2.3. Polytopes
The combinatorial automorphism group is generally difficult to compute since the faces are precisely what interests us. Again considering the extreme rays $\{\mathbb{R}v_{i}\}_{1\leq i\leq m}$ for some representative vectors $v_{1},\ldots,v_{m}$, the restricted linear automorphism group $\operatorname{RLin}_{.}(P)$ is the group of permutations $\sigma\in\operatorname{Sym}(n)$ such that there exists a matrix $B\in\operatorname{GL}_{n}(\mathbb{R})$ for which $Bv_{i}=v_{\sigma(i)}$ for all $1\leq i\leq m$. Note that the obtained group can depend on the precise representatives $v_{i}$. However, we do always have that $\operatorname{RLin}_{.}(P)\subset\operatorname{Comb}(P)$. Furthermore, this group is efficiently computable in practice by reducing it to a graph automorphism problem. We refer to [BDSP+14] for more theoretical and computational details. For our purposes it is sufficient to only obtain the facet orbits up to some symmetry group $G\subset\operatorname{Comb}(P)$.
@@ 3. Ryshkov’s polyhedra and perfect forms
There are two viewpoints on perfect forms that have been considered, which are essentially dual to each other. The original one, introduced by Voronoi [Vor08], considers perfect domains, and is extensively treated by Martinet [Mar03]. The other point of view, which considers Ryshkov polyhedra, was introduced in [Rys70] and is detailed further in [Sch09b]. Both viewpoints are considered together in [Opg01] in a generalized setting. The Ryshkov polyhedron gives a more intuitive view with respect to the lattice packing problem, so we focus on this notion here. The idea is to restrict the space of lattice bases (up to isometry), to those with a (squared) first minimum of at least some $\lambda>0$. It turns out that under the Frobenius inner product, this results in a convex body.
Definition 4 (Ryshkov polyhedron [Rys70]).
@@
An example of a Ryshkov polyhedron in dimension $2$ is shown in fig. 2. Note that Ryshkov polyhedra are technically not polyhedra according to our definition, as they are an infinite intersection of halfspaces. However they are locally finite [Sch09a, Theorem 3.1]: the intersection of $\mathcal{P}^{d}_{\lambda}$ with any other polytope $P$, is again a polytope, i.e., locally they behave the same as a polyhedron. In particular, we can therefore still naturally talk about vertices and facets of $\mathcal{P}^{d}_{\lambda}$, even if there are an infinite number of them. With this parallel, note that the facets of $\mathcal{P}^{d}_{\lambda}$ correspond precisely to all primitive integer vectors $\pm x\in\mathbb{Z}^{d}$ up to sign.
Figure 2. On the left, part of Ryshkov polyhedron inside the cone of $2$-dimensional positive definite quadratic forms. Its vertices are perfect forms, such as the red and green vertices on the right. The group $\operatorname{GL}_{d}(\mathbb{Z})$ acts as a symmetry group on the Ryshkov polyhedron. The red vertices are neighbouring perfect forms of the green vertex, indicating an exploration step of Voronoi’s algorithm.
@@
The function $Q\mapsto\det(Q)^{1/d}$ was already shown to be concave in $\mathcal{S}_{>0}^{d}$ by Minkowski [Min05] (see [vW18, Theorem 6] for an elementary proof). It is even strictly concave except under scaling, in particular it is strictly concave on the boundary of $P_{\lambda}^{d}$. Because $\mathcal{P}_{\lambda}^{d}$ is locally finite  [Sch09a, Theorem 3.1] and convex the minimum of the function over $\mathcal{P}_{\lambda}^{d}$ is thus attained at the vertices. ∎
@@
Theorem 3 (Number of perfect forms [Bac18, vW20]).
@@ 4. Voronoi’s algorithm
The classical Voronoi algorithm [Vor08] is a graph traversal algorithm that enumerates all perfect forms. That algorithm is explained in the books [Sch09a, Mar03] and articles [DSSV07, Sch09b].
@@ 4.1. The algorithm
Algorithm 1 Voronoi’s algorithm [Vor08]
Dimension $d\geq 1$. A complete list LLof non-similar perfect forms of dimension $d$.
[…]
2:Compute $\operatorname{Min}(Q)$.
3:Enumerate the extreme rays $R_{1},\ldots,R_{k}$ of the cone
$\mathcal{P}(Q)=\{Q^{\prime}\in\mathcal{S}^{d}:\langle Q^{\prime},xx^{t}\rangle\geq 0\text{ for all }x\in\operatorname{Min}(Q)\}.$
[…]
@@
Computing the extreme rays of $\mathcal{P}(Q)$ from the known facets, or equivalently, computing the facets of $\operatorname{Vor}(Q)=\mathcal{P}(Q)^{\circ}$ from the known extreme rays, is precisely a dual description problem. For certain cases this problem is relatively easy, for example when $|\operatorname{Min}(Q)|=d(d+1)$ there are precisely $\frac{1}{2}d(d+1)$ linearly independent facets equal to the dimension of the space $\mathcal{S}^{d}$. In this case $\mathcal{P}(Q)$ is a simplicial cone with precisely $\frac{1}{2}d(d+1)$ extreme rays, and computing these can be done relatively quickly. Informally, we call these cases of low incidence. We can efficiently solve these low-incidence cases with specialized software like LRS [Avi15], CDD [Fuk15] and PPL [BHZ08].
@@
Given a perfect form $Q$ and an extreme ray $R$ of $\mathcal{P}(Q)$ the algorithm needs to compute a scalar $\alpha>0$ such that $Q^{\prime}=Q+\alpha R$ is a perfect form with the same minimum. Note that for all $x\in\operatorname{Min}(Q)$ we have $x^{t}Rx\geq 0$ by definition of $R$ being an extreme ray of $\mathcal{P}(Q)$. Those minimal vectors that correspond to the facets adjacent to the extreme ray satisfy $x^{t}Rx=0$ and therefore will also end up in $\operatorname{Min}(Q^{\prime})$. We need to find the minimal $\alpha>0$ such that $Q+\alpha R$ lies on a new facet, i.e., the minimal $\alpha>0$ such that $\operatorname{Min}(Q^{\prime})$ contains a new vector not contained in $\operatorname{Min}(Q)$. One can approach this by a binary search, i.e., for such a new vector $y\in\operatorname{Min}(Q^{\prime})\setminus\operatorname{Min}(Q)$ we must have $y^{t}Ry<0$ and thus for $\alpha^{\prime}>\alpha$ we have that either $Q+\alpha^{\prime}R$ is not positive definite, or that $\lambda_{1}(Q+\alpha^{\prime}R)<\lambda_{1}(Q)$. In the latter case we can also compute any $y\in\operatorname{Min}(Q+\alpha^{\prime}R)<\lambda_{1}(Q)$ and decrease $\alpha^{\prime}$ to the point that it satisfies $y^{t}(Q+\alpha^{\prime}R)y=\lambda_{1}(Q)$. For $\alpha^{\prime}<\alpha$ we have $\operatorname{Min}(Q+\alpha^{\prime}R)\subset\operatorname{Min}(Q)$ and thus we can also detect this. This binary search process can be further optimized in practice using LLL and a fast implementation for the enumeration of short vectors. For a more extensive treatment of this step, both from a practical and an asymptotic point of view we refer to  [vW18, Chapter 5].

### v2

@@ 1. Introduction
The sphere packing problem is a classical and notoriously hard problem. It asks how to pack equal $n$-dimensional balls in $\mathbb{R}^{n}$ such that their density is maximized. The sphere packing problem has only been solved in dimensions $1,2,3,8$ and $24$, the last two of which only recently [viazovska2017sphere, cohn2017sphere]. Johannes Kepler conjectured already in the 17th century that in dimension $3$ no packing could be denser than the face-centered cubic or hexagonal close packings. A testament of the hardness of this problem is given by the fact that Kepler’s conjecture stayed unproven for almost $400$ years. Until $1998$, when Thomas Hales announced a computational proof to the sphere packing problem in dimension $3$ [hales2005kepler] which was later formalized with proof checking software in [hales2017formal].
@@
And indeed, this restriction allows to solve the problem in all dimensions $1$ to $8$ and $24$. Contrary to Kepler’s conjecture the lattice packing problem in dimension $3$ was already solved by Gauss in $1831$ [gauss1831besprechung]. Dimensions $4$ and $5$ followed in $1877$ by Korkine and Zolotarev’s reduction theory [korkine1877formes], and in $1935$ Blichfeldt used the same theory to determine the optimal lattice packings in dimensions $6,7$ and $8$ [blichfeldt1935minimum]. The famous Leech lattice was shown to be optimal in dimension $24$ by Cohn and Kumar [cohn2009leechoptimality] using linear programming bounds, a technique which also played an important role in solving the general sphere packing problem in dimensions $8$ and $24$. The optimality in all these dimensions was shown by strong theoretical arguments and limited computations. However, the lattice packing problem in dimensions $9$ and beyond (except for $24$) so far seems to evade any of the theoretical approaches that worked for the other dimensions. In particular, all known density upper bounds are relatively far from the best known (lattice) packings in these dimensions.
Almost $100$ years before the computational proof of the Kepler conjecture, Voronoi, ahead of his time, already gave a computational approach to solve the lattice packing problem in any fixed dimension [VoronoiI]. Voronoi showed that any optimal lattice packing must be a perfect lattice, and that there are only a finite number of non-similar perfect lattices in any given dimension. Furthermore, he presented an algorithm that provably enumerates all perfect lattices, thereby giving a computational way to solve the lattice packing problem.
@@
2 1 Lagrange, 1770 [de1770demonstration]
3 1 Gauss, 1831 [gauss1831besprechung]
4 2 Korkine & Zolotarev, 1877 [korkine1877formes]
5 3 Korkine & Zolotarev, 1877 [korkine1877formes]
6 7 Barnes, 1957 [barnes1957perfect]
7 33 Jaquet-Chiffelle, 1993 [jaquet1993enumeration]
@@
9 $\geq$ 500 000 Vallentin, 2005 [PerfectDim8]
$\geq$ 23 000 000 van Woerden, 2018 [vanWoerdenMasterThesis]
@@ 9 $\geq$ 500 000 Vallentin, 2005 [PerfectDim8]
So far, this computational approach to solve the lattice packing problem, has always fallen behind the theoretical results, mainly due to its high complexity. For example, the enumeration of perfect lattices in dimension $6$ was done by Barnes in $1957$ [barnes1957perfect], $22$ years after Blichfeldt’s proof. And dimensions $7$ and $8$ followed, only in $1993$ by Jaquet [jaquet1993enumeration] and in $2007$ by Dutour Sikirić, Schürmann and Vallentin [PerfectDim8] respectively. Dimension $9$ was also attempted, and in [PerfectDim8] and [AnzinEnumeration] about 500 000 perfect lattices were found. In [vanWoerdenMasterThesis] some additional algorithmic improvements lead to a total of 23 000 000 perfect lattices to be found. Unfortunately, even with such a large number there was still no end in sight during the enumeration, and provably finding all cases requires a significant amount of extra computation, making it unclear if dimension $9$ would be feasible.
@@
Given that $\gamma_{8}=\gamma_{9}=2$ this also shows that the Hermite constant does not have to be strictly increasing in the dimension. Furthermore, additional data obtained by our computation allows us to prove precisely how many shortest vectors a lattice in dimension $9$ can have.
@@
We aim to make the data and implementations used in this work public. The full classification of perfect forms in dimension $9$ is available split into 64 binary netCDF files at [DutourSikiric2025Complete]. Their total size is roughly 150GB. We wrote C++ and python utilities to easily access this data once downloaded [^1] Available at https://github.com/MathieuDutSik/AccessPerfectDim9 and https://github.com/WvanWoerden/AccessPerfectDim9Py respectively.. The dual description under symmetry program has been developed publicly at [PolyhedralCpp]. Our parallel implementation of Voronoi’s algorithm, including the canonical form implementation, will be made public in due course.
@@ Acknowledgements
The authors would like to thank Achill Schürmann and Jacques Martinet for insightful discussions and suggestions that were beneficial to this work. The authors would also like to thank Matthias Schymura and Gregory Minton for useful feedback on the first draft of this work.
@@ 2.1. Lattices and lattice packings
Example 1.
As an example let use consider the best lattice packing $\mathsf{E}_{8}$ in dimension $8$ and the (until this work) conjectured best lattice packing $\mathsf{\Lambda}_{9}$ in dimension $9$. The $\mathsf{E}_{8}$ root lattice has a relative simple description as follows
$\mathsf{E}_{8}:=\{x\in\mathbb{Z}^{8}\cup(\mathbb{Z}+\tfrac{1}{2})^{8}:\sum_{i=1}^{8}x_{i}\equiv 0\bmod{2}\}.$
It has a determinant of $1$, a first minimum of $\sqrt{2}$ and $240$ minimal vectors, leading to a density of $\pi^{4}/(2^{4}\cdot 4!)\approx 0.25367$ which was shown to be optimal by Blichfeldt [blichfeldt1935minimum]. The laminated lattice $\mathsf{\Lambda}_{9}$ can be constructed by stacking the $\mathsf{E}_{8}$ root lattice. The unit vector $e_{1}=(1,0,0,0,0,0,0,0)$ is a deep hole of $\mathsf{E}_{8}$ on which we can stack another copy of $\mathsf{E}_{8}$ by using the gluing vector $e_{1}+e_{9}$ of length $\sqrt{2}$. We then obtain (one version of) the laminated lattice as follows
$\mathsf{\Lambda}_{9}:=\{(x~|~0)\in\mathbb{R}^{9}~|~x\in\mathsf{E}_{8}\}+(e_{1}+e_{9})\cdot\mathbb{Z}.$
This lattice has determinant $1$, a first minimum of $\sqrt{2}$ and $272$ minimal vectors, leading to a density of $(4!\cdot 2^{9/2}\cdot\pi^{4})/9!\approx 0.14577$. Typically we scale $\mathsf{\Lambda}_{9}$ by $\sqrt{2}$ to make it integral.
@@ 2.3. Polytopes
The combinatorial automorphism group is generally difficult to compute since the faces are precisely what interests us. Again considering the extreme rays $\{\mathbb{R}v_{i}\}_{1\leq i\leq m}$ for some representative vectors $v_{1},\ldots,v_{m}$, the restricted linear automorphism group $\operatorname{RLin}_{.}(P)$ is the group of permutations $\sigma\in\operatorname{Sym}(n)$ such that there exists a matrix $B\in\operatorname{GL}_{n}(\mathbb{R})$ for which $Bv_{i}=v_{\sigma(i)}$ for all $1\leq i\leq m$. Note that the obtained group can depend on the precise representatives $v_{i}$. However, we do always have that $\operatorname{RLin}_{.}(P)\subset\operatorname{Comb}(P)$. Furthermore, this group is efficiently computable in practice by reducing it to a graph automorphism problem. We refer to [GroupPolytopeLMS] for more theoretical and computational details. For our purposes it is sufficient to only obtain the facet orbits up to some symmetry group $G\subset\operatorname{Comb}(P)$.
@@ 3. Ryshkov’s polyhedra and perfect forms
There are two viewpoints on perfect forms that have been considered, which are essentially dual to each other. The original one, introduced by Voronoi [VoronoiI], considers perfect domains, and is extensively treated by Martinet [martinet]. The other point of view, which considers Ryshkov polyhedra, was introduced in [ryshkov1970polyhedron] and is detailed further in [achill-enumerating]. Both viewpoints are considered together in [Opgenorth2001] in a generalized setting. The Ryshkov polyhedron gives a more intuitive view with respect to the lattice packing problem, so we focus on this notion here. The idea is to restrict the space of lattice bases (up to isometry), to those with a (squared) first minimum of at least some $\lambda>0$. It turns out that under the Frobenius inner product, this results in a convex body.
Definition 4 (Ryshkov polyhedron [ryshkov1970polyhedron]).
@@
An example of a Ryshkov polyhedron in dimension $2$ is shown in fig. 2. Note that Ryshkov polyhedra are technically not polyhedra according to our definition, as they are an infinite intersection of halfspaces. However they are locally finite [bookschurmann, Theorem 3.1]: the intersection of $\mathcal{P}^{d}_{\lambda}$ with any other polytope $P$, is again a polytope, i.e., locally they behave the same as a polyhedron. In particular, we can therefore still naturally talk about vertices and facets of $\mathcal{P}^{d}_{\lambda}$, even if there are an infinite number of them. With this parallel, note that the facets of $\mathcal{P}^{d}_{\lambda}$ correspond precisely to all primitive integer vectors $\pm x\in\mathbb{Z}^{d}$ up to sign.
Figure 2. On the left, part of Ryshkov polyhedron inside the cone of 22-dimensional positive definite quadratic forms. Its vertices are perfect forms, such as the red and green vertices on the right. The group $\operatorname{GL}_{d}(\mathbb{Z})$ acts as a symmetry group on the Ryshkov polyhedron. The red vertices are neighbouring perfect forms of the green vertex, indicating an exploration step of Voronoi’s algorithm.
@@
The function $Q\mapsto\det(Q)^{1/d}$ was already shown to be concave in $\mathcal{S}_{>0}^{d}$ by Minkowski [minkowski1905diskontinuitatsbereich] (see [vanWoerdenMasterThesis, Theorem 6] for an elementary proof). It is even strictly concave except under scaling, in particular it is strictly concave on the boundary of $P_{\lambda}^{d}$. Because $\mathcal{P}_{\lambda}^{d}$ is locally finite  [bookschurmann, Theorem 3.1] and convex the minimum of the function over $\mathcal{P}_{\lambda}^{d}$ is thus attained at the vertices. ∎
@@
Theorem 3 (Number of perfect forms [bacher2018number, van2020upper]).
@@ 4. Voronoi’s algorithm
The classical Voronoi algorithm [VoronoiI] is a graph traversal algorithm that enumerates all perfect forms. That algorithm is explained in the books [bookschurmann, martinet] and articles [PerfectDim8, achill-enumerating].
@@ 4.1. The algorithm
Algorithm 1 Voronoi’s algorithm [VoronoiI]
1: Dimension $d\geq 1$.
2: A complete list $L$ of non-similar perfect forms of dimension $d$.
3: $L\leftarrow\{Q_{\mathsf{A}_{d}}\}$ $\triangleright$ perfect form corresponding to root lattice $\mathsf{A}_{d}$.
4: while $\exists$ untreated $Q\in L$ do
5:   Compute $\operatorname{Min}(Q)$.
6:   Enumerate the extreme rays $R_{1},\ldots,R_{k}$ of the cone $\mathcal{P}(Q)=\{Q^{\prime}\in\mathcal{S}^{d}:\langle Q^{\prime},xx^{t}\rangle\geq 0\text{ for all }x\in\operatorname{Min}(Q)\}.$
7:   for all extreme rays $R_{i}$ of $\mathcal{P}(Q)$ do
8:    Determine neighbouring perfect form $Q_{i}=Q+\alpha_{i}R_{i}$. $\triangleright$ $\alpha_{i}>0$
9:    if $Q_{i}$ is not equivalent to any form in $L$ then
10:      $L\leftarrow L\cup\{Q_{i}\}$.
11:   Mark $Q\in L$ as treated.
12: output $L$
@@
Computing the extreme rays of $\mathcal{P}(Q)$ from the known facets, or equivalently, computing the facets of $\operatorname{Vor}(Q)=\mathcal{P}(Q)^{\circ}$ from the known extreme rays, is precisely a dual description problem. For certain cases this problem is relatively easy, for example when $|\operatorname{Min}(Q)|=d(d+1)$ there are precisely $\frac{1}{2}d(d+1)$ linearly independent facets equal to the dimension of the space $\mathcal{S}^{d}$. In this case $\mathcal{P}(Q)$ is a simplicial cone with precisely $\frac{1}{2}d(d+1)$ extreme rays, and computing these can be done relatively quickly. Informally, we call these cases of low incidence. We can efficiently solve these low-incidence cases with specialized software like LRS [lrs], CDD [cdd] and PPL [BagnaraHZ08SCP].
@@
Given a perfect form $Q$ and an extreme ray $R$ of $\mathcal{P}(Q)$ the algorithm needs to compute a scalar $\alpha>0$ such that $Q^{\prime}=Q+\alpha R$ is a perfect form with the same minimum. Note that for all $x\in\operatorname{Min}(Q)$ we have $x^{t}Rx\geq 0$ by definition of $R$ being an extreme ray of $\mathcal{P}(Q)$. Those minimal vectors that correspond to the facets adjacent to the extreme ray satisfy $x^{t}Rx=0$ and therefore will also end up in $\operatorname{Min}(Q^{\prime})$. We need to find the minimal $\alpha>0$ such that $Q+\alpha R$ lies on a new facet, i.e., the minimal $\alpha>0$ such that $\operatorname{Min}(Q^{\prime})$ contains a new vector not contained in $\operatorname{Min}(Q)$. One can approach this by a binary search, i.e., for such a new vector $y\in\operatorname{Min}(Q^{\prime})\setminus\operatorname{Min}(Q)$ we must have $y^{t}Ry<0$ and thus for $\alpha^{\prime}>\alpha$ we have that either $Q+\alpha^{\prime}R$ is not positive definite, or that $\lambda_{1}(Q+\alpha^{\prime}R)<\lambda_{1}(Q)$. In the latter case we can also compute any $y\in\operatorname{Min}(Q+\alpha^{\prime}R)<\lambda_{1}(Q)$ and decrease $\alpha^{\prime}$ to the point that it satisfies $y^{t}(Q+\alpha^{\prime}R)y=\lambda_{1}(Q)$. For $\alpha^{\prime}<\alpha$ we have $\operatorname{Min}(Q+\alpha^{\prime}R)\subset\operatorname{Min}(Q)$ and thus we can also detect this. This binary search process can be further optimized in practice using LLL and a fast implementation for the enumeration of short vectors. For a more extensive treatment of this step, both from a practical and an asymptotic point of view we refer to  [vanWoerdenMasterThesis, Chapter 5].